OpenStudy (anonymous):

how do you find the max height for y = -5x^2 + 24t

OpenStudy (amoodarya):

y = -5x^2 + 24t o ry = -5x^2 + 24x ?

OpenStudy (anonymous):

oh sorry my bad it's y = -5t^2 + 24t

OpenStudy (anonymous):

what, dont you do completing the square?

OpenStudy (amoodarya):

ok lets do it factor -5

OpenStudy (anonymous):

so -5(t^2 + 24t)

OpenStudy (amoodarya):

no -5(t^2-24/5t) -5(t^2-4.8t)

OpenStudy (amoodarya):

-5(t^2-4.8t +(2.4^2)-(2.4^2))

OpenStudy (amoodarya):

-5(t^2-4.8t +(2.4^2)-(2.4^2)) =-5[(t-2.4)^2-2.4^2} max obtain by t=2.4